The volume strength of 1.5 N H2 O2 is
a. 4.8
b. 8.4
c. 3.0
d. 8.0

Grade:10

3 Answers

Jitender Pal
9 years ago
Sol. Volume strength of H2O2 = Normality × 5.6 = 1.5 × 5.6 = 8.4 V
Sarthak Gupta
5 years ago
Strenth =normality * eq weight.. =1.5*17=25.5 68g of h
Yash Chourasiya
3 years ago
Dear Student

We know that equivalent weight of H2O2 is 17
and strength of H2O2=Normality x Equivalent weight
= 1.5x 17 = 25.5
Since 68 grams of H2O2 produces 22.4 litres oxygen at NTP
Therefore 25.5 grams of H2O2 will produce = 22.4/68 x 25.5 = 8.4 litre of oxygen
Thus, volume strength of given H2O2 solution is 8.4

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya